# Group Testing with General Correlation Using Hypergraphs

### v1

@@ 6.3 When $\mu$ is a Lower Bound for Achieving Zero Probability of Error
Graph $G$ has this property that for every node $v$ and every realization of $V\setminus v$, there is a positive chance that $v$ is negative and there is a positive chance that $v$ is negative. In other words, for every $e\in E$ and $u\notin e$, there is an edge $e^{\prime}=\{v\}\cup e$ where $e^{\prime}\in E$.
@@
We claim that any testing algorithm on the graph in Example 6.6 with zero probability of error needs at least $\mu$ tests.
@@
First, we prove that every positive node $v$ should be presented in a test in order to achieve a zero probability error. If not, even knowing the states of $V\setminus v$, there is a positive chance that $v$ is positive or negative, which contradicts the assumption of zero error. Second, we prove that each positive node should be at least in a test where every other node in the test is negative. Suppose not, and for every test for a node $v$ that is positive there is another positive node. Then by changing the state of $v$ from positive to negative, the results of the tests remain the same. However having $v$ as a negative node corresponds to a valid edge by the assumption, contradicting the zero error assumption. This shows that for any graph with the realized edge $e^{*}$, at least $|e^{*}|$ tests are needed, meaning that on average, at least $\mu$ tests are needed. ∎
6.4 Random Graphs and Regimes of Optimality
Consider $d$-regular random graphs, where each subset $S\subseteq V$ of size $|S|=d$ is included in the graph with probability $r$, independent of the rest, to make the graph $G_{r}=(V,E_{r})$. Then $G_{r}$ is revealed with a distribution $\mathcal{D}$ defined on the edges. We analyze two regimes of interest by examining the upper and lower bounds in each case. The first regime is when $r$ is small and the graph is sparse, and the other is when $r$ is large and the graph is dense. The following shows that when $r$ is large, there is a distribution $\mathcal{D}$ over the edges $E_{r}$ so that $\mu$ is a lower bound.
@@ 6.4 Random Graphs and Regimes of Optimality
Suppose graph $G_{r}=(V,E_{r})$ is given with the parameter $r\gg(\frac{d}{n})^{1/d}$. Then there is a distribution $\mathcal{D}$ on $E_{r}$ so that any algorithm with a constant probability of success needs $O(\mu)=O(d)$ number of tests on average.
@@
Suppose that $r\gg(\frac{d}{n})^{1/d}$. We partition the nodes into groups of size $d+1$, so there are $\frac{n}{d+1}$ groups. The probability that all the edges in one group are sampled is $r^{d+1}$, so with probability $(1-r^{d+1})^{\frac{n}{d+1}}$, no such group has all its edges sampled. Because $(1-r^{d+1})^{\frac{n}{d+1}}\simeq e^{-r^{d+1}{\frac{n}{d+1}}}$, if $r\gg(\frac{d}{n})^{1/d}$, this probability goes to zero. Hence, with high probability, there are $d+1$ nodes where every subset of size $d$ amongst them is a sampled edge. Call such a subset $S$ and all the the edges inside $e_{1},e_{2},\ldots,e_{d+1}$, where $\forall i:e_{i}\in E_{r}$ and $e_{i}\subset S$ and $|S|=d+1$. Now define the distribution $\mathcal{D}$ such that it only puts weights on $e_{1},e_{2},\ldots,e_{d+1}$, uniformly at random, i.e. $\forall i:\mathcal{D}(e_{i})=\frac{1}{d+1}$. Following the line of argument in Example 6.1, we conclude that any algorithm with a constant probability of success needs at least $O(\mu)=O(d)$ number of tests. ∎
Note that Algorithm 1 performs $O(H(X)+\mu)=O(H(X)+d)$ in the above regime by Theorem 4.8, and we already know that $H(X)$ is a lower bound. The above proposition shows that $O(d)$ is also necessary, proving that Algorithm 1 is order-wise optimal in this regime.
@@
Suppose graph $G_{r}=(E_{r},V)$ is given with the parameter $r\ll(\frac{d}{n})^{O(d)}$. Then Algorithm 1 finds the infected set using $O(H(X))$ tests on average.
@@
We prove that when $r\ll(\frac{d}{n})^{O(d)}$, with high probability there is no subset of nodes of size at most a constant of $d$, let’s say $2d$ where more than a constant number of edges amongs them are sampled. If this is true, then when the Algorithm 1 gets to the Line 1 (recall that this is the line that the algorithm performs individual testing), there is only a constant number of edges remaining and it can test each of the edges instead of each of the nodes. Note that testing one edge here is possible using only one test. This is because the graph is $d-$regular and by testing $V\setminus e$, we can check if $e$ is the target edge or not.
The probability that more than $c^{\prime}$ edges, for a constant $c^{\prime}$, are sampled among $2d$ nodes is less than $r^{c^{\prime}}{2d\choose c^{\prime}}<(rd/(c^{\prime}/2))^{c^{\prime}}$. There are (${n\choose 2d}$ candidates for picking up a subset of size $2d$, and (${n\choose 2d}<(\frac{n}{2d})^{2d}$. Hence by union bound, the probability that there is a subset of size $2d$ with more than $c^{\prime}$ edges is at most $(\frac{n}{2d})^{2d}(rd/(c^{\prime}/2))^{c^{\prime}}$. When $r\ll(\frac{d}{n})^{O(d)}$, this goes to zero and hence with high probability, fewer than $c$ edges exist when Algorithm 1 gets to Line 1 and performs at most $c^{\prime}$ tests.
So with high probability, the total number of tests is $O(H(X))$ in the first phase of the algorithm on average, and $c^{\prime}$ tests in the second phase. So the total number of tests is $O(H(X)+c^{\prime})=O(H(X))$ on average, which shows that Algorithm 1 is order-wise optimal.
The above proposition shows that the analysis provided in Theorem 4.10 is worst-case scenario and even though we have a term $\mu=d$ in the upper bound of the theorem, Algorithm 1 is order-wise optimal.

### v2

@@ 6.3 When $\mu$ is a Lower Bound for Achieving Zero Probability of Error
Graph $G$ has this property that for every node $v$ and every realization of states of $V\setminus v$, there is a positive chance that $v$ is infected and there is a negative chance that $v$ is not infected. Thus, for every subset of nodes there is a positive probability that only the nodes in it are positive and the nodes outside are negative, hence, every subset of nodes is a target edge with a positive probability and therefore belongs in $E$.
@@
Any algorithm on Example 6.8 with zero probability of error needs at least $\mu$ tests in expectation.
@@
Let $e^{*}$ be the target edge. Then, for zero probability of error, $e^{*}$ has to be detected as the target edge. Thus, every node in $e^{*}$ has to be detected as positive. Thus, every node in $e^{*}$ has to be tested at least once alone or along with other nodes as its state can not be definitively inferred even if states of all other nodes in $V$ are known, from the condition in the Example. We next argue that every node in $e^{*}$ has to be tested at least once alone or in a group in which no other node in $e^{*}$ is present. This is because all other nodes in $e^{*}$ are positive, and if the considered node in $e^{*}$ is tested only in company of other positive nodes, the outcome of all such tests would be positive regardless of the state of the considered node. Thus, since each node in $e^{*}$ is tested at least once without any other node in $e^{*}$ being tested in the same test, the number of tests must be at least $|e^{*}|.$ The result follows because $E|e^{*}|=\mu.$
6.4 Random Hypergraphs and Regimes of Optimality
Consider $d$-regular random hypergraphs, (ie, where all edges have $d$ nodes, hence $\mu=d$), and each subset $S\subseteq V$ of size $|S|=d$ is included in it with probability $r$, independent of the rest, to make the graph $G_{r}=(V,E_{r})$. We analyze two regimes of interest by examining the upper and lower bounds in each case. The first regime is when $r$ is large and the graph is dense (eg, $r$ is $\omega((d/n)^{1/d})$), the other is when $r$ is small (eg, $O((d/n)^{\Theta(d)})$) and the graph is sparse.
@@ 6.4 Random Hypergraphs and Regimes of Optimality
Suppose graph $G_{r}=(V,E_{r})$ is given with the parameter $r\gg(\frac{d+1}{n})^{1/(d+1)}$. Then there is a distribution $\mathcal{D}$ on $E_{r}$ so that any algorithm with a constant probability of success needs $\Omega(\mu)=\Omega(d)$ number of tests on average.
The Proposition and its proof are based on the following insight. Since $r$ is large and the hypergraph $G_{r}$ is dense, enough number of subsets of size $d$ from $V$ make it to $E_{r}$, such that with a high probability there exists at least one set $S\subset V$ of size $d+1$ such that all subsets of size $d$ of $S$ constitute edges in $E_{r}.$ Now, Example 6.3, Claim 6.4 and its proof, show that for such a graph, at least $\Omega(d)$ tests are needed to identify the target edge with a constant probability of success, if the probability distribution for the target edge is uniform over the edges of size $d$ in $S$.
@@
Suppose that $r\gg(\frac{d+1}{n})^{1/(d+1)}$. We partition the nodes into subsets of size $d+1$, so there are $\frac{n}{d+1}$ sets. Each such subset has $d+1$ edges of size $d$ each. The probability that all the edges of size $d$ in a given subset are included in $G_{r}$ is $r^{d+1}$, so with probability $(1-r^{d+1})^{\frac{n}{d+1}}$, no such subset has all its edges sampled. Note that $(1-r^{d+1})^{\frac{n}{d+1}}\leq e^{-r^{d+1}{\frac{n}{d+1}}}.$ Denote $e^{-r^{d+1}{\frac{n}{d+1}}}$ as $\delta$. Hence, with a probability $1-\delta$ there is at least one subset of $d+1$ nodes, say $S$, where every subset of size $d$ in it is an edge in $G_{r}$. We first show that $\delta$ is a small positive number. Let $r^{\prime}=r^{d+1}\frac{n}{d+1}.$ Then, $\delta=e^{-r^{d+1}{\frac{n}{d+1}}}=e^{-r^{\prime}}$, and $\frac{r}{(\frac{d+1}{n})^{1/(d+1)}}=r^{\prime}.$ Since $r\gg(\frac{d+1}{n})^{1/(d+1)}$, $r^{\prime}\gg 1.$ Thus, $e^{-r^{\prime}}$ is a small positive number.
Consider that there is such a subset $S$ as described in the first paragraph. Call the edges inside $S$ of size $d$, $e_{1},e_{2},\ldots,e_{d+1}$, where $\forall i:e_{i}\in E_{r}$ and $e_{i}\subset S$ and $|S|=d+1$. Now define the distribution $\mathcal{D}$ such that it only puts weights on $e_{1},e_{2},\ldots,e_{d+1}$, uniformly at random, i.e. $\forall i:\mathcal{D}(e_{i})=\frac{1}{d+1}$. Following the line of argument in proof of Claim 6.4 for Example 6.3, we conclude that any algorithm with a constant probability of success needs $\Omega(d)$ number of tests in expectation. Note that this holds if such an $S$ exists which happens with probability $1-\delta$. If such a $S$ does not exist, even allowing that 00 tests are needed to ensure constant probability of success, the overall expected number of tests is $(1-\delta)\Omega(d)$, which is also $\Omega(d).$ Also, $\mu=d$, so every such algorithm needs $\Omega(\mu)$ tests in expectation. ∎
Note that Algorithm 1 performs $O(H(X)+\mu)=O(H(X)+d)$ tests in the above regime by Theorem 4.9, and we already know that $H(X)$ is a lower bound (Theorem 2.2). The above proposition shows that $d$ is also a lower bound, in the asymptotic order sense, proving that Algorithm 1 is order-wise optimal in this regime.
We now consider the sparse regime with small $r:$
@@
Suppose the random hypergraph $G_{r}=(E_{r},V)$ is given with the parameter $r$ as $O((d/n)^{c_{r}d})$, where $c_{r}=\frac{2}{c^{\prime}(1-2c)},$ $c$ is the constant parameter defined in Algorithm 1, and $c^{\prime}$ is any constant positive integer. Then, there is a modified version of Algorithm 1 that finds the infected set using $O(H(X))$ tests on average when $d$ is $o(n)$, and $H(X)$ is $\Omega(1)$.
Proposition 6.11 and its proof are based on the following insight. Algorithm 1 uses $O(H(X)+d)$ tests in expectation for a $d$-regular hypergraph because it tests nodes of $S$ individually in line 1 and $S$ has $O(\mu)$ nodes in expectation just before line 1 and for a $d$-regular Hypergraph $\mu=d$. But for a sparse $d$-regular random hypergraph, with a high probability $S$ has only a constant number of edges of $E_{r}.$ This is because any set of cardinality $O(d)$ has $O((n/d)^{O(d)})$ subsets of cardinality $d$, and if $r$ is $O((d/n)^{O(d)})$, then only a constant number of these subsets are in $E_{r}$ in expectation. Thus, $|E(S)|$ is constant in expectation. But not all sets of cardinality $O(d)$ has the target edge in it. Before its line 1, Algorithm 1 identifies a set $S$ of cardinality $O(d)$ which has the target edge in it. Thus, instead of testing nodes in $S$ individually, the edges in $E(S)$ can be tested to determine which of them is the target edge. This would need a constant number of tests in expectation in line 1 instead of $O(d)$ tests therein, if there exists a way to directly test an edge. Note that in general it is not possible to conclude that an edge is a target edge by testing all its nodes together as a group, because the outcome will be positive even if one node in it is positive; thus one group test can not determine if all nodes in it are positive. But this can be accomplished by testing the complement of each edge in $E(S)$ utilizing the fact that $G_{r}$ is a $d-$regular graph. Thus, $d$ in the upper bound on the number of tests executed by Algorithm 1 can be replaced by a constant and the Proposition follows as $H(X)$ is at least a constant.
@@
Consider a modified version of Algorithm 1 catering specifically to regular hypergraphs (ie, where all edges have equal sizes). In line 1, the modified version first checks if the number of remaining edges in $S$ is fewer than $|S|$. If it holds, the algorithm tests each $S\setminus e$ for each remaining edge $e\in E(S)$ and outputs $e$ as the target edge the first time one such test is negative. If the condition does not hold, then it tests nodes in $S$ individually and outputs the positive nodes as the target edge. Note that the target edge has to be one of the remaining edges in $S$ as edges are removed only when it is established that they can not be the target edge. Thus since all positive nodes are in the target edge, the test of $S\setminus e$ will be negative for the target edge $e.$ If $S\setminus e$ is negative, either $e$ or a proper subset thereof is the target edge. But since all remaining edges are of equal size, the target edge can not be a proper subset of any remaining edge. Hence, the modified algorithm concludes that $e$ is the target edge. Note that the modified algorithm needs at most $\min(|E(S)|,|S|)$ tests in the last step.
We now specifically consider the $d$-regular random hypergraph $G_{r}.$ Recall that the original Algorithm 1 performs at most $\frac{\tilde{\mu}}{1-2c}$ tests in line 1 (Lemma 4.8), where $\tilde{\mu}$ is the expected number of infected nodes at the start of line 1. Since each edge consists of $d$ nodes, so does the target edge, and thus, there are exactly $d$ infected nodes in the network, so $\tilde{\mu}=d.$ Also, the original Algorithm 1 performs exactly $|S|$ tests in line 1. So $|S|\leq\frac{d}{1-2c}.$ Since the modified algorithm needs at most $\min(|E(S)|,|S|)$ tests in the last step, the expected number of tests in the last step is $O(d)$. So if $d=O(1)$, the expected number of tests in line 1 is also $O(1).$ We prove that even when $d$ is $\omega(1)$, the expected number of tests in line 1 is $O(1).$ Then the expected number of tests by the modified algorithm is 1) $O(H(X))$ until before line 1 (from Corollary 4.6 since the modified and original versions differ only in line 1) 2) at most $O(1)$ in line 1. Since $H(X)=\Omega(1)$, the total number of tests is $O(H(X))$ on average. The Proposition follows.
In the rest of the proof we assume that $d$ is $\omega(1)$, and using the fact that every edge has $d$ nodes, we prove that with probability $1-\epsilon_{n}$, where $\lim_{n\rightarrow\infty}\epsilon_{n}=0$, the number of edges of $E_{r}$ in any subset of $V$ of cardinality at most $\frac{d}{1-2c}$ is less than any given constant positive integer $c^{\prime}$ when $r$ is small and $d$ is $o(n).$ Thus, with probability $1-\epsilon_{n}$, $|E(S)|<c^{\prime}$, when the original Algorithm 1 gets to line 1, and the same holds for the modified version as they differ only in line 1. Thus, with probability $1-\epsilon_{n}$, $c^{\prime}$ is an upper bound on the number of tests the modified algorithm needs in this last step. Since the number of tests in this last step is always upper bounded by $|S|$, and $|S|\leq\frac{d}{1-2c}$, if there are $c^{\prime}$ or more edges in $S$, which happens with probability at most $\epsilon_{n}$, we consider the upper bound to be $\frac{d}{1-2c}$ instead. Thus, the expected number of tests in this final step is at most $c^{\prime}+\epsilon_{n}(d/(1-2c)).$ We will prove that $\lim_{n\rightarrow\infty}\epsilon_{n}d=0$ using the fact that $d$ is $o(n).$ Thus, since $c$ is a constant, in the limit that $n\rightarrow\infty$, the upper bound for the expected number of tests in line 1 is $c^{\prime}$, which would complete the proof of the Proposition.
We now prove that for any constant positive integer $c^{\prime}$, with probability $1-\epsilon_{n}$, where $\lim_{n\rightarrow\infty}\epsilon_{n}=0$, the number of edges in any subset of $V$ of cardinality at most $\frac{d}{1-2c}$ is less than $c^{\prime}$. First note that if there does not exist $S$ of a certain cardinality with at least $c^{\prime}$ edges, there does not exist $S$ with a lower cardinality with at least $c^{\prime}$ edges, because a set is always a subset of some other set with higher cardinality. Thus, the event that there does not exist a set with at least $c^{\prime}$ edges, is a subset of the event that there does not exist a set of higher cardinality with at least $c^{\prime}$ edges. Thus, the event that there does not exist a set with cardinality less than or equal to $\frac{d}{1-2c}$ with at least $c^{\prime}$ edges is the event that there does not exist a set with cardinality equal to $\lfloor\frac{d}{1-2c}\rfloor$ with at least $c^{\prime}$ edges. We therefore upper bound the probability of the latter. Let $\gamma=1/(1-2c).$ First, there are (${\lfloor\gamma d\rfloor\choose d}$ subsets of size $d$ in a given set of cardinality $\lfloor\gamma d\rfloor.$ Also, (${\lfloor\gamma d\rfloor\choose d}\leq(\frac{e\lfloor\gamma d\rfloor}{d})^{d}\leq(e\gamma)^{d}$ as (${n\choose k}\leq(\frac{en}{k})^{k}$, for each $n,k\leq n.$ Now, the probability that at least $c^{\prime}$ such subsets are in $E_{r}$ is less than $r^{c^{\prime}}{{\lfloor\gamma d\rfloor\choose d}\choose c^{\prime}}\leq(re(\gamma e)^{d}/c^{\prime})^{c^{\prime}}$. There are (${n\choose\lfloor\gamma d\rfloor}$ candidates for picking up a subset of size $\lfloor\gamma d\rfloor$ from $V$, and
(${n\choose(\lfloor\gamma d\rfloor)}\leq(\frac{en}{\lfloor\gamma d\rfloor})^{\lfloor\gamma d\rfloor}\leq(\frac{en}{d})^{\lfloor\gamma d\rfloor}\leq(\frac{en}{d})^{\gamma d}$
, since $d\leq n$, $d$ is $\omega(1)$. Hence by union bound, the probability that there exists $S$ of size $\lfloor\gamma d\rfloor$ with at least $c^{\prime}$ edges is at most
$(\frac{en}{d})^{\gamma d}(re(\gamma e)^{d}/c^{\prime})^{c^{\prime}}=(\gamma e)^{c^{\prime}d}(\frac{en}{d})^{\gamma d}(re/c^{\prime})^{c^{\prime}}=(e/(1-2c))^{c^{\prime}d}(\frac{en}{d})^{d/(1-2c)}(re/c^{\prime})^{c^{\prime}},$
replacing $\gamma$ with $1/(1-2c)$. Let $\kappa=(e/(1-2c))^{c^{\prime}(1-2c)}$, and $\nu=(e/c^{\prime})^{c^{\prime}}$. Then, we get $(e/(1-2c))^{c^{\prime}d}(\frac{en}{d})^{d/(1-2c)}(re/c^{\prime})^{c^{\prime}}=\nu(\kappa\frac{n}{d})^{d/(1-2c)}(r)^{c^{\prime}}$. When $r\leq(\frac{d}{\kappa n})^{\frac{2d}{c^{\prime}(1-2c)}}$, the right hand side is at most $\nu(\frac{d}{\kappa n})^{\frac{d}{c^{\prime}(1-2c)}}=\epsilon_{n}$ where $\epsilon_{n}$ goes to zero as $n\rightarrow\infty$ since $d$ is $o(n)$ and $d$ is $\omega(1)$. Note that the upper bound we considered for $r$ is $O((d/n)^{c_{r}d})$, where $c_{r}=\frac{2}{c^{\prime}(1-2c)}$.
We now prove that $\lim_{n\rightarrow\infty}\epsilon_{n}d=0$. We have
$\epsilon_{n}d=\nu(\frac{d}{\kappa n})^{\frac{d}{c^{\prime}(1-2c)}}d=\nu(\frac{d^{1+\frac{c^{\prime}(1-2c)}{d}}}{\kappa n})^{\frac{d}{c^{\prime}(1-2c)}}.$
As $d^{1/d}<2$, $d^{\frac{c^{\prime}(1-2c)}{d}}<2^{\frac{c^{\prime}(1-2c)}{d}}<2^{c^{\prime}},$ we get $\epsilon_{n}d<\nu(\frac{2^{c^{\prime}}d}{\kappa n})^{\frac{d}{c^{\prime}(1-2c)}}$. Since $d$ is $o(n),$ $\lim_{n\rightarrow\infty}\epsilon_{n}d=0$.
The parameter $c_{r}$ controls the degree of sparsity of the hypergraph considered in Proposition 6.11. Specifically, $\lim_{n\rightarrow\infty}\mathbb{E}[|E_{r}|]$ can range from infinity to 00, depending on the value of $c_{r}$. To see this note that the number of subsets of $V$ of cardinality $d$, ie, ($n\choose d$, is at least $(n/d)^{d}.$ The expected number of edges in $E_{r}$ is therefore at least $r(n/d)^{d}$, which can be $\Theta((n/d)^{(1-c_{r})d})$ if $r$ is $\Theta((d/n)^{c_{r}d})$. If $0<c_{r}<1,$ $(n/d)^{(1-c_{r})d}\rightarrow\infty$ as $n\rightarrow\infty$, since $d$ is $o(n).$ Thus $\lim_{n\rightarrow\infty}\mathbb{E}[|E_{r}|]=\infty.$ Since $c_{r}=\frac{2}{c^{\prime}(1-2c)},$ $20<c<1/$ 2, for all $c^{\prime}>2/(1-2c)$, $10<c_{r}<$ 1 . Recall that $|E(S)|$ is constant in expectation for the $S$ identified by Algorithm 1 and its modification before their respective final steps, even though $\lim_{n\rightarrow\infty}\mathbb{E}[|E_{r}|]=\infty.$ This can be ensured only because $S$ has $O(d)$ (instead of $n$) nodes, and for each $S\subset V$, with $O(d)$ nodes, $|E(S)|$ is constant in expectation, and before their final steps Algorithm 1 and its modification identify an $S$ with $O(d)$ nodes which has the target edge, and therefore in the last step the algorithm can solely focus on $S$ and the edges therein. But, if the hypergraph becomes even sparser, ie, $c_{r}$ higher, leading to lower $r$, then $|E_{r}|$ itself becomes a constant or goes to zero in expectation. For example, if $c_{r}>1$, (ie, $1<c^{\prime}<2/(1-2c)$), then since there are $O((n/d)^{d})$ subsets of size $d$ of $V$, then $\mathbb{E}[E_{r}]=O((n/d)^{d}(d/n)^{c_{r}d})=O((d/n)^{(c_{r}-1)d})\rightarrow 0$, as $n\rightarrow\infty$ (since $d$ is $o(n)$, $c_{r}>1$). In other words, the hypergraph is empty, and no node is infected. No algorithm is meaningful in this case.
Note that the proof of Proposition 6.11 shows that if $r$ is $O((d/n)^{c_{r}d})$, $c_{r}=\frac{2}{c^{\prime}(1-2c)},$ $c$ is the constant parameter defined in Algorithm 1, then $|E(S)|<c^{\prime}$ with a high probability for each $S\subset V$, with $O(d)$ nodes. For large $c^{\prime}$, ie, $c^{\prime}>2/(1-2c)$, the requirement is lax and can be met even when the $d$-regular random hypergraph has a large edge set, ie, $\lim_{n\rightarrow\infty}\mathbb{E}[|E_{r}|]=\infty.$ Thus, $r$ does not need to be very low in this case, nor $G_{r}$ extremely sparse, which corresponds to $10<c_{r}<$ 1 . But for small $c^{\prime}$, ie, $c^{\prime}<2/(1-2c)$, the requirement is stringent and can be met only when the $d$-regular random hypergraph has a vanishing edge set, ie, $\lim_{n\rightarrow\infty}\mathbb{E}[|E_{r}|]=0.$ Thus $G_{r}$ needs to be extremely sparse, actually nearly empty, to start with, ie, $r$ very low, which happens for high $c_{r}$, ie, $c_{r}>1.$
Finally, if $d$ is $\alpha n$, instead of $o(n)$, then again the total number of subsets of size $d$ of $V$ is small enough so that the hypergraph is empty regardless of the value of $c_{r}.$ We delve into the details and justify various assumptions made in Proposition 6.11 next.
Remark 6.12.
For the assumption in Proposition 6.11 that $d$ is $o(n)$, if we assume otherwise, i.e., $d$ is $\Theta(n)$, or $d=\alpha n$ for a constant $\alpha\leq 1$, then the random hypergraph becomes so sparse that no node is infected in the asymptotic limit on $n$. This can be seen as follows. The total number of subsets of size $d$ is (${n\choose d}={n\choose\alpha n}\leq(e/\alpha)^{n}$. Hence if $r$ is $O((d/n)^{c_{r}d})$ $r$ is $O((\frac{\alpha n}{n})^{(c_{r}\alpha n)})=O((\alpha^{c_{r}\alpha})^{n})$. Since $0<\alpha\leq 1$, as long as $c_{r}\geq 0$, $\alpha^{c_{r}\alpha}\leq 1<e/\alpha$, then $\mathbb{E}(|E_{r}|)$ is at most $c_{1}^{n}$ for a constant $c_{1}<1$. Thus $\mathbb{E}(|E_{r}|)$ goes to 0 as $n\rightarrow\infty$. Since $Pr(|E_{r}|\geq 1)\leq\mathbb{E}(|E_{r}|)$, $Pr(|E_{r}|\geq 1)$ also goes to 0 as $n\rightarrow\infty$. Thus, the random hypergraph becomes empty in the asymptotic limit on $n$.
The assumption $H(X)=\Omega(1)$ is natural, as if $H(X)=o(1)$, it means that the distribution for target edge is concentrated on a single edge and if the probability of error is larger than the probability that this edge is not the target edge, one can output this edge without any test. In other words, the assumption $H(X)=\Omega(1)$, limits considerations to instances where at least one test is needed to identify the target edge.
Proposition 6.11 and Theorem 2.2 together imply that the modified version of Algorithm 1 is order-wise optimal for sparse $d$-regular random hypergraphs when $d$ is $o(n)$, and the distribution for the target edge is not excessively concentrated on one edge. Also, the analysis provided in Theorem 4.11 is the worst-case scenario and a slight modification of the algorithm renders the $\mu=d$ term in the upper bound of the theorem as redundant in many cases.